# WorldForge: Unlocking Emergent 3D/4D Generation in Video Diffusion Model via Training-Free Guidance

### v1

@@ 4 Experiments
In this section, we present a comprehensive evaluation of our proposed training-free framework. We first outline the implementation details in Sec. 4.1. Subsequently, we demonstrate the performance of our method on 3D scene generation and 4D trajectory control in Sec. 4.2 and Sec. 4.3, respectively. Finally, we conduct a series of ablation studies in Sec. 4.4 to validate the effectiveness of each component of our approach.
@@ 4.1 Implementation Details
Our framework is a training-free, inference-time optimization method that steers pre-trained video diffusion models for precise camera control without additional training or fine-tuning. It introduces no significant computational overhead beyond the base model’s inference requirements.
@@
Our pipeline follows a warp-and-repaint design. For warping, we test several depth estimation models, such as VGGT (Wang et al., 2025b), UniDepth (Piccinelli et al., 2024), Mega-SaM (Li et al., 2025), and DepthCrafter (Hu et al., 2025). The method adapts well to all, benefiting from the strong world priors of the underlying video model.
Test Datasets and Metrics. For single-view 3D scene generation, we adopt the widely used LLFF (Mildenhall et al., 2019), Tanks and Temples (Knapitsch et al., 2017), and MipNeRF 360 (Barron et al., 2022) datasets, selecting half of the scenes from each for evaluation. We also test on diverse internet, real-world, and AI-generated images. Perceptual quality is assessed using FID (Heusel et al., 2017) and $\mathrm{CLIP}_{\mathrm{sim}}$ (Radford et al., 2021) similarity. For 4D trajectory control, we compare with SOTA methods on their respective benchmarks, which include challenging real-world video clips with varied camera trajectories.
Performance is measured with FVD (Unterthiner et al., 2018) for temporal quality and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ for source–target consistency. In addition, we assess camera trajectory accuracy using three standard metrics: Absolute Trajectory Error (ATE), which captures the global alignment between estimated and reference paths; Relative Pose Error – Translation (RPE-T), which reflects short-term consistency of translational motion; and Relative Pose Error – Rotation (RPE-R), which quantifies local orientation accuracy. Together, these metrics provide a comprehensive evaluation of both global fidelity and local smoothness of camera trajectories.
4.2 3D Scene Generation
Figure 3: Qualitative comparison of novel view synthesis from a single input image. The first row shows the input frame and its depth-based warped views, where disoccluded regions appear as holes. Compared to existing SOTA methods, our approach produces more consistent scene content under novel viewpoints, with improved image detail, trajectory accuracy, and structural plausibility.
We compare our method with SOTA approaches for novel view synthesis, which can be grouped into two categories. The first includes methods requiring model-specific training or fine-tuning, such as ViewCrafter (Yu et al., 2024c), which trains a VDM on iteratively refined point clouds; TrajectoryCrafter (Yu et al., 2025), which learns motion control via dedicated training; TrajectoryAttention (Xiao et al., 2025), which injects motion through a specialized attention module; and See3D(Ma et al., 2025a), a video-based model for scene reconstruction. The second category comprises training-free methods, such as NVS-Solver (You et al., 2025), which modulates a pre-trained model using warped views, and ViewExtrapolator (Liu et al., 2024), which employs an annealed guidance strategy to improve trajectory adherence.
Table 1: Quantitative results on static 3D and dynamic 4D benchmarks. We evaluate on public datasets such as LLFF and MipNeRF360, as well as Internet images and videos. For static scenes we report $\mathrm{CLIP}_{\mathrm{sim}}$ and FID, and for dynamic scenes $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ and FVD. Metric details are provided in Appendix B. Our method consistently surpasses SOTA baselines in generation quality.
Static Dynamic
FID $\downarrow$ $\mathrm{CLIP}_{\mathrm{sim}}\uparrow$ FVD $\downarrow$ $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}\uparrow$
@@
96.08 0.948 93.17 0.938
Table 2: Quantitative comparison of camera-trajectory accuracy on static and dynamic scenes. We evaluate against SOTA methods using ATE, RPE-T, and RPE-R, lower is better. Across both settings, our method achieves the best or second-best results on all metrics. Metric definitions and computation details are provided in Appendix B.
Static Dynamic
ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
See3D (Ma et al., 2025a)
0.091 0.089 0.250 - - -
ViewCrafter (Yu et al., 2024c)
0.236 0.315 0.728 - - -
ViewExtrapolator (Liu et al., 2024)
0.183 0.260 0.882 1.040 1.208 4.750
TrajectoryAttention (Xiao et al., 2025)
0.159 0.238 0.532 0.605 1.238 3.560
TrajectoryCrafter (Yu et al., 2025)
0.090 0.152 0.267 0.431 1.078 8.950
NVS-Solver (You et al., 2025)
0.224 0.268 1.056 - - -
WorldForge (Ours)
0.077 0.086 0.221 0.527 0.826 2.690
[…]
Frames Resolution
Inference Time
Base Video
Model
Training-
Free
See3D (Ma et al., 2025a)
25 $576\times 1024$ 1.7 Custom
ViewCrafter (Yu et al., 2024c)
25 $576\times 1024$ 1.8 Custom
ViewExtrapolator (Liu et al., 2024)
25 $576\times 1024$ 1.6 SVD
TrajectoryAttention (Xiao et al., 2025)
25 $76\times 1024 5.5$ SVD
TrajectoryCrafter (Yu et al., 2025)
25 $384\times 672$ 1.7 CogVideoX
NVS-Solver (You et al., 2025)
25 $576\times 1024$ 9.3 SVD
ReCamMaster (Bai et al., 2025a)
81 $480\times 832$ 14.6 Wan 2.1 T2V
WorldForge (Ours, 720P)
25 $1280\times 720$ 17.3 Wan 2.1 I2V
WorldForge (Ours, 480P)
25 $480\times 832$ 6.8 Wan 2.1 I2V
WorldForge (Ours, on SVD)
25 $576\times 1024$ 1.3 SVD
We evaluate all methods under consistent settings with the same input image, camera trajectory, depth, and target scenes across LLFF (Mildenhall et al., 2019), MipNeRF-360 (Barron et al., 2022), and Tanks-and-Temples (Knapitsch et al., 2017). Results (Figs.3, Table1) show that our inference-time guidance consistently outperforms both training-based and training-free baselines. Moreover, we further assess human-centric scenes, which are more sensitive to visual consistency and content quality. As shown in Fig. 4, prior methods often fail in such cases. In contrast, our approach, supported by the proposed guidance strategy, preserves model priors and faithfully follows the target trajectory. This offers an effective paradigm for reconciling controllability with generalization and high fidelity. To independently validate the effect of our guidance strategy, we further evaluate on the lighter U-Net-based SVD model (Blattmann et al., 2023); even in this setting, our method attains superior visual quality.
Notably, our framework can synthesize high-quality, photorealistic, and structurally consistent $360^{\circ}$ views from a single input without relying on panoramic intermediates (Fig. 5). This capability extends to complex outdoor and open-world scenes, pushing the limits of single-view novel view synthesis.
4.3 4D Trajectory Control
For dynamic trajectory control, we compare against leading video-to-video models. These include ReCamMaster (Bai et al., 2025a), which employs a sophisticated conditioning mechanism trained on large-scale synthetic data, TrajectoryCrafter (Yu et al., 2025), with its dual-stream diffusion architecture, TrajectoryAttention (Xiao et al., 2025), which processes motion through an additional attention module, and ViewExtrapolator (Liu et al., 2024).
Following the protocols of prior works, we evaluate on diverse and challenging video clips, including arcs, dolly zooms, and composite paths. Our training-free method consistently delivers higher visual fidelity, better trajectory alignment, and more coherent scene completion. As shown in Table 1 and Fig. 6, it matches or surpasses state-of-the-art methods that require costly training, demonstrating the strength of leveraging latent 3D priors in existing models. Notably, our approach excels at plausibly reconstructing unseen regions, while other methods often produce distortions or implausible content.
Figure 6: Comparison of 4D trajectory-controlled re-rendering. Baselines often produce implausible artifacts (e.g., flattened faces, floating heads), reflecting limited use of pretrained priors. Our inference-time guidance leverages these latent world priors to re-render realistic, high-quality content along the target trajectory. We compare against state-of-the-art baselines under identical inputs; for ReCamMaster (text-controlled), parameters are adjusted to match the target path.
Beyond benchmarks, the framework supports various post-production tasks. It can stabilize videos by smoothing camera motion and control camera paths to enable localized super-resolution or outpainting. In addition, by specifying masked regions, it supports creative video content edits such as object addition or removal, subject replacement, and try-on effects (see Fig. 7). These capabilities highlight its versatility for real-world video re-rendering.
@@ 4.4 Ablation Experiments
We conducted ablation studies to assess the contribution of each component in our framework.
@@
Component Analysis. We examined the effects of removing Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG). As shown in Fig. 8 without IRR, trajectory guidance cannot be injected during inference, and the model fails to follow the target path. Removing FLF, i.e., not separating motion and appearance channels, damages the model’s priors and produces unnatural results. Without DSG, noise from warped trajectories propagates into the generation, causing low quality and artifacts. The full model achieves the best results, confirming that these components work synergistically to ensure robust and precise control.
Base Model. We replaced the Wan2.1 (Wan et al., 2025) model with the U-Net-based SVD (Blattmann et al., 2023) model to verify model-agnosticism and transferability (See in Fig. 9). Fair comparisons with other SVD-based methods show that our approach integrates seamlessly into pre-trained SVD. Although limited by SVD’s weaker priors, our method still reaches state-of-the-art performance within the same model. This demonstrates that our guidance is architecture-independent and suggests even greater potential with stronger future base models.
Depth Estimation Model. As shown in Fig. 10 We tested VGGT (Wang et al., 2025b), Mega-SaM (Li et al., 2025), UniDepth (Piccinelli et al., 2024) and DepthCrafter (Hu et al., 2025) for depth-based warping. Each shows distinct strengths and weaknesses, yet our method remains effective across all. Thanks to the generative model’s strong 3D prior, many warping artifacts (e.g., tearing, stretching) are automatically corrected during refinement. This robustness enables plug-and-play integration with various SOTA depth estimation techniques without sacrificing performance.
Figure 10: Depth estimation models ablation. Our method leverages the inherent world knowledge of VDMs to correct errors and fill missing regions even under challenging inputs (left). This strong self-correction ability ensures broad compatibility with different depth estimators (right). Despite variations or noise in depth-based warping, it reliably compensates through learned priors and produces realistic, high-quality results.
5 Conclusion and Limitation
@@
We present WorldForge, a training-free framework for trajectory-controllable generation in both static 3D scenes and dynamic 4D scenes. Our method tackles the persistent challenge in controllable video generation of balancing high visual quality, strong generalization, and precise controllability. At its core is a unified guidance strategy—Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG)—that operates entirely at inference time. By decoupling motion from appearance features and correcting trajectory drift caused by structural noise, it injects fine-grained predefined trajectory constraints while preserving the valuable prior world knowledge embedded in the base model. Extensive experiments demonstrate state-of-the-art performance on both 3D and 4D trajectory generation tasks, offering a new paradigm for exploring spatial intelligence and emergent world models in large-scale generative systems.
@@
While our framework can repair many distortions introduced by depth-based warping, it may fail under extremely poor depth estimations (e.g., completely flattened subjects or severe foreground–background entanglement). Moreover, due to the global nature of our guidance, control over small objects or fine details remains limited. Future work will explore integrating fine-grained control mechanisms and applying our approach to more powerful generative models.
@@ B.1 Static Scene Evaluation
Fréchet Inception Distance (FID). FID (Heusel et al., 2017) measures image generation quality by comparing the distribution of real and generated images in the Inception-V3 feature space. We use an ImageNet-pretrained Inception-V3 model and extract 2048-dimensional features from the pool3 layer. The FID score is computed as:
@@ B.2 Dynamic Scene Evaluation
Fréchet Video Distance (FVD). FVD (Unterthiner et al., 2018) extends FID to the video domain by measuring distribution differences in spatio-temporal feature space. For computational efficiency, we implement a simplified version that treats video frame sequences as image collections and applies Inception-V3 frame-wise feature extraction. The FVD score follows the same Fréchet distance formula as FID but operates on video frame features.
@@ B.3 Camera Trajectory Evaluation
Absolute Trajectory Error (ATE). ATE measures the global consistency between estimated and reference camera trajectories by computing Euclidean distances between corresponding camera positions. The ATE for each timestamp is computed as:
$\text{ATE}_{i}=\|\mathbf{t}_{\text{ref},i}-\mathbf{t}_{\text{est},i}\|_{2}$ (28)
where $\mathbf{t}_{\text{ref},i}$ and $\mathbf{t}_{\text{est},i}$ are the 3D position vectors of the reference and estimated trajectories at timestamp $i$, respectively. The root mean square error over all $n$ trajectory points is:
$\text{ATE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\text{ATE}_{i}^{2}}$ (29)
Relative Pose Error - Translation (RPE-T). RPE-T evaluates the local accuracy of camera translation between consecutive frames, reflecting short-term trajectory consistency. For consecutive frames, the relative translation error is:
$\text{RPE-T}_{i}=\|\Delta\mathbf{t}_{\text{ref},i}-\Delta\mathbf{t}_{\text{est},i}\|_{2}$ (30)
where $\Delta\mathbf{t}_{\text{ref},i}=\mathbf{t}_{\text{ref},i+1}-\mathbf{t}_{\text{ref},i}$ and $\Delta\mathbf{t}_{\text{est},i}=\mathbf{t}_{\text{est},i+1}-\mathbf{t}_{\text{est},i}$. The RMSE over all n1n-1 consecutive frame pairs is:
$\text{RPE-T}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-T}_{i}^{2}}$ (31)
Relative Pose Error - Rotation (RPE-R). RPE-R assesses the accuracy of camera orientation changes between consecutive frames. For rotation matrices $\mathbf{R}_{\text{ref},i}$ and $\mathbf{R}_{\text{est},i}$, the angular difference in degrees is computed as:
$\text{RPE-R}_{i}=\arccos\left(\frac{\text{trace}(\Delta\mathbf{R}_{\text{ref},i}^{T}\Delta\mathbf{R}_{\text{est},i})-1}{2}\right)\cdot\frac{180}{\pi}$ (32)
where $\Delta\mathbf{R}_{\text{ref},i}=\mathbf{R}_{\text{ref},i+1}\mathbf{R}_{\text{ref},i}^{T}$ and $\Delta\mathbf{R}_{\text{est},i}=\mathbf{R}_{\text{est},i+1}\mathbf{R}_{\text{est},i}^{T}$. The RMSE over all consecutive frame pairs is:
$\text{RPE-R}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-R}_{i}^{2}}$ (33)

### v2

@@ 4 Experiments
In this section, we present a comprehensive evaluation of our proposed training-free framework. We first outline the implementation details in Sec. 4.1. Subsequently, we demonstrate the performance of our method on 3D scene generation and 4D trajectory control in Sec. 4.2. Finally, we conduct a series of ablation studies in Sec. 4.3 to validate the effectiveness of each component of our approach. More results and details are provided in Appendix E.
Figure 3: Qualitative comparison for 3D novel view synthesis and 4D trajectory-controlled re-rendering. (Left) For 3D scene generation from a single image, our method produces more consistent and plausible content compared to SOTA baselines. (Right) For 4D re-rendering, our approach leverages model priors to avoid common artifacts like floating heads and flattened faces, yielding more realistic results. Overall, our training-free guidance demonstrates superior performance in both static and dynamic scenes. Zoom in for more details.
@@ 4.1 Implementation Details
Our framework is a training-free method that steers pre-trained VDMs for precise camera control, with no additional training or fine-tuning. Compared with pre-trained VDMs, our method incurs no training cost, while the inference time increases by approximately 40–50%, mainly due to the IRR module (see Appendix E.1 for detailed efficiency analysis).
@@
Our pipeline follows a warp-and-repaint design. For warping, we test several depth estimation models, including VGGT (Wang et al., 2025b), UniDepth (Piccinelli et al., 2024), Mega-SaM (Li et al., 2025), and DepthCrafter (Hu et al., 2025). The method adapts well across these choices, benefiting from the VDM’s strong world priors; see Appendix E.3 for depth-estimator ablations.
Test Datasets and Metrics. For single-view 3D scene generation, we use data from LLFF (Mildenhall et al., 2019), Tanks and Temples (Knapitsch et al., 2017), MipNeRF 360 (Barron et al., 2022), and diverse internet, real-world, and AI-generated images. We report FID (Heusel et al., 2017) and $\mathrm{CLIP}_{\mathrm{sim}}$ (Radford et al., 2021). For 4D trajectory control, we use challenging real-world videos, reporting FVD (Unterthiner et al., 2018) and $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$. For both tasks, trajectory accuracy is measured with Absolute Trajectory Error (ATE), Relative Pose Error—Translation (RPE-T), and Relative Pose Error—Rotation (RPE-R). Full metric definitions are in Appendix B.
4.2 3D and 4D Trajectory-Controlled Generation
We compare our method against state-of-the-art baselines on both 3D static scene generation and 4D dynamic video control. For 3D novel view synthesis, we evaluate against both training-based (Yu et al., 2024c; 2025; Xiao et al., 2025; Ma et al., 2025a) and training-free (You et al., 2025; Liu et al., 2024) methods. For 4D trajectory control, baselines include ReCamMaster (Bai et al., 2025a) and others (Yu et al., 2025; Xiao et al., 2025; Liu et al., 2024).
Under identical evaluation settings, our training-free method consistently achieves superior results. On 3D static scenes, it outperforms both training-based and training-free baselines on public datasets (Fig. 3, Table 1). On 4D clips with challenging camera paths (e.g., arcs, dolly zooms), it yields higher visual fidelity, tighter trajectory alignment, and more coherent scene completion, matching or surpassing costly training-based approaches. In both settings, our method plausibly reconstructs unseen regions where baselines often produce distortions.
Our approach particularly excels in difficult cases. It handles human-centric scenes, which require high consistency, and can synthesize photorealistic and structurally consistent $360^{\circ}$ views from a single input (Appendix E.5). By preserving model priors, it strikes a strong balance between controllability and fidelity. We tested our method on a lighter SVD model (Blattmann et al., 2023) and still achieved high visual quality, confirming its strong performance across different model scales (Appendix E.2).
Beyond benchmarks, the framework serves as a versatile tool for video post-production. It can stabilize videos by smoothing camera motion, control paths for localized super-resolution or outpainting, and perform masked edits like object addition/removal or subject try-on effects. These capabilities highlight its utility for real-world video re-rendering (more results in Appendix E.4).
Table 1: Quantitative comparison with SOTA methods on 3D static and 4D dynamic scenes, using public and internet data. We evaluate generation quality (FID, $\mathrm{CLIP}_{\mathrm{sim}}$ for static; FVD, $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ for dynamic) and trajectory accuracy (ATE, RPE-T, RPE-R). All methods use official code with identical inputs. $\uparrow$: Higher is better, $\downarrow$: Lower is better. Our method achieves the best or second-best results on all metrics. Metric details are in Appendix B.
Generation Quality Trajectory Accuracy
Static Dynamic Static Dynamic
FID $\downarrow$ $\mathrm{CLIP}_{\mathrm{sim}}\uparrow$ FVD $\downarrow$ $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
@@
96.08 0.948 93.17 0.938 0.077 0.086 0.221 0.527 0.826 2.690
@@ 4.3 Ablation Experiments
We ablate each module and design choice in our framework.
@@
Component Analysis. We remove IRR, FLF, and DSG in turn (Fig. 4). Removing IRR disables trajectory guidance at inference time, resulting in failure to follow the target path. Without FLF, i.e., lacking motion/appearance separation, model priors become entangled, leading to unnatural outputs. Removing DSG introduces noise from warped trajectories into the generation process, causing artifacts and degrading visual quality. The complete model yields the best results, showing that all components are essential and work synergistically to enable robust and precise control.
Video Model and Depth Model. We replace Wan 2.1 (Wan et al., 2025) with the U-Net–based SVD (Blattmann et al., 2023) to test model-agnosticism. It consistently achieved strong results, demonstrating its effectiveness across a range of model architectures(Appendix E.2). Similarly, our method remains robust regardless of the warping model. We experimented with different depth estimators (Wang et al., 2025b; Li et al., 2025; Piccinelli et al., 2024; Hu et al., 2025) and found that the VDM’s strong 3D priors effectively mitigate many warping artifacts. This allows for a plug-and-play integration with various depth models (Appendix E.3).
5 Conclusion
@@
We present WorldForge, a training-free framework for trajectory-controllable generation in static 3D and dynamic 4D scenes. Our method effectively balances visual quality, generalization, and precise control in video synthesis. At its heart is a unified, inference-time guidance strategy—comprising Intra-Step Recursive Refinement (IRR), Flow-Gated Latent Fusion (FLF), and Dual-Path Self-Corrective Guidance (DSG). By decoupling motion from appearance and correcting trajectory drift, our framework injects fine-grained control while preserving the rich world priors of the base model. Extensive experiments show state-of-the-art performance on both 3D and 4D generation tasks, offering a new path for exploring spatial intelligence in large-scale generative systems.
@@
While our framework corrects many warping-induced distortions, it can fail with severely inaccurate depth estimations (e.g., completely flattened subjects or severe foreground–background entanglement). Furthermore, the global nature of our guidance offers limited control over small objects and fine details. Future work will focus on integrating fine-grained control mechanisms and applying our method to more powerful generative models.
@@ B.1 Static Scene Evaluation
Fréchet Inception Distance (FID). FID (Heusel et al., 2017) measures image generation quality by comparing the distribution of real and generated images in the Inception-V3 feature space. We use an ImageNet-pretrained Inception-V3 (Szegedy et al., 2016) model and extract 2048-dimensional features from the pool3 layer. The FID score is computed as:
@@ B.2 Dynamic Scene Evaluation
Fréchet Video Distance (FVD). FVD (Unterthiner et al., 2018) measures distributional differences between real and generated video using pretrained spatio-temporal features. We use an I3D (Inflated 3D ConvNet) pretrained on Kinetics (Carreira & Zisserman, 2017) and extract 1024-D features from the global average pooling layer for each video clip. Following FID, we compute the Fréchet distance between the Gaussian fits of real and generated I3D features:
$\text{FVD}=\|\mu_{r}-\mu_{g}\|_{2}^{2}+\mathrm{Tr}\!\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}\right),$ (21)
with $\mu_{r},\mu_{g}$ and $\Sigma_{r},\Sigma_{g}$ estimated over clip-level I3D features.
@@ B.3 Camera Trajectory Evaluation
Absolute Trajectory Error (ATE). Before evaluation, we align the estimated trajectory to the reference by a global Sim3 transform (scale, rotation, translation). Let the aligned pose components be $\tilde{\mathbf{t}}_{\text{est},i}$ and $\tilde{\mathbf{R}}_{\text{est},i}$. ATE measures global consistency by the Euclidean distance between corresponding camera positions:
$\text{ATE}_{i}=\big\|\mathbf{t}_{\text{ref},i}-\tilde{\mathbf{t}}_{\text{est},i}\big\|_{2},\qquad\text{ATE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\text{ATE}_{i}^{2}}.$ (23)
Relative Pose Error — Translation (RPE-T). RPE-T evaluates local translation accuracy between consecutive frames. Define relative motions via poses (index gap $\Delta\!=\!1$):
$\Delta\mathbf{T}_{\text{ref},i}=\mathbf{T}_{\text{ref},i}^{-1}\mathbf{T}_{\text{ref},i+1},\qquad\Delta\mathbf{T}_{\text{est},i}=\tilde{\mathbf{T}}_{\text{est},i}^{-1}\tilde{\mathbf{T}}_{\text{est},i+1}.$ (24)
Let $\Delta\mathbf{t}_{\text{ref},i}$ and $\Delta\mathbf{t}_{\text{est},i}$ be the translation parts of these relative transforms. The per-step error and RMSE are:
$\text{RPE-T}_{i}=\big\|\Delta\mathbf{t}_{\text{ref},i}-\Delta\mathbf{t}_{\text{est},i}\big\|_{2},\qquad\text{RPE-T}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-T}_{i}^{2}}.$ (25)
Relative Pose Error — Rotation (RPE-R). RPE-R assesses the accuracy of orientation changes between consecutive frames. Let the relative rotations be
$\Delta\mathbf{R}_{\text{ref},i}=\mathbf{R}_{\text{ref},i}^{-1}\mathbf{R}_{\text{ref},i+1},\qquad\Delta\mathbf{R}_{\text{est},i}=\tilde{\mathbf{R}}_{\text{est},i}^{-1}\tilde{\mathbf{R}}_{\text{est},i+1}.$ (26)
The per-step angular error (degrees) and RMSE are:
$\text{RPE-R}_{i}=\arccos\!\left(\frac{\mathrm{trace}\!\big(\Delta\mathbf{R}_{\text{ref},i}^{\top}\Delta\mathbf{R}_{\text{est},i}\big)-1}{2}\right)\cdot\frac{180}{\pi},\qquad\text{RPE-R}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n-1}\text{RPE-R}_{i}^{2}}.$ (27)